# Enhancing Gain in Non-Hermitian Photonic Crystals with Lossy Topological Defects

### v1

@@ Abstract
Figure 3: (a) schematics of the baseline periodic and line-defected 2D non-Hermitian system. The baseline periodic system is shown on the right with 51 unit cells of length a. Each unit cell is surrounded by air and composed of a lossless rod of $\varepsilon_{a}=2$ and gain rod of $\varepsilon_{b}=2-0.1i$ with radii $0.2a$ and $0.1a$ respectively. While the photonic crystal is finite in the x direction, it is infinitely periodic in the y direction from imposing floquet periodic boundary conditions. The defected case is shown on the right which contains a lossy line defect composed of rods with $\varepsilon_{d}$ and radius $r_{d}=0.3*a$. The defect is surrounded by 25 unit cells of the baseline periodic system on the left and right extents. (b) Gain vs frequency plots of the baseline periodic crystal compared against the lossy line defected case at loss values of 1, 1.5, and 1.75.
@@
We now consider an analogous two-dimensional non-Hermitian system. The baseline periodic 2D non-Hermitian photonic crystal contains of 51 unit cells that repeat in the $x$ direction of rods surrounded by air similar to the 1D baseline periodic structure (Fig. 3(a)). Within each unit cell, there is one lossless rod of $\varepsilon_{a}=2$ and a rod with gain of $\varepsilon=2-0.1i$. The lossless rod has radius $r_{a}=0.2a$ while the gain rod has radius $r_{b}=0.1a$ where $a$ is the unit cell length. It is also important to note that, though the photonic crystal is finite in the x-direction, Floquet periodic boundary conditions are imposed to make the system infinitely periodic in the $y$-direction as well. In the defected case, a geometry similar to the 1D point-defected structure is used where 25 unit cells surround either side of a lossy or lossless defect rod defined by a permittivity $\varepsilon_{d}=\varepsilon^{\prime}_{d}+\varepsilon^{\prime\prime}_{d}i$ and radius $r_{d}=0.3a$, resulting in the same $C_{2}$ symmetry as the 1D system. Since Floquet periodic boundary conditions were placed on the boundaries in the $y$-direction, a line defect is created instead of a point defect in this system. 4a.
@@
Figure 4: (a), (b) Complex reflection phase as quantified by $Arg(TE^{TE}_{11})$ vs $Im(\varepsilon_{d})$ at a $\varepsilon^{\prime}_{d}=2$ and $Re(\varepsilon_{d})$ at 0 loss of the line defect respectively. Increasing $Im(\varepsilon_{d})$ leads to a transition from a winding number of +1 to 0, while the complex reflection phase remains continuous with changing $Re(\varepsilon_{d})$.
@@
As in the 1D system, increasing the loss in the point defect up to a critical value leads to multiple orders of magnitude in gain enhancement Fig. 3(b). This behavior can be understood by examining the integer winding of the complex reflection phase in the frequency vs $\varepsilon_{d}$ parameter space. Modulating the loss, as shown in Fig. 4(a), within the defect induces a topological phase transition around a critical loss value of $\varepsilon^{\prime\prime}_{d}=1.5$ (while maintaining $\varepsilon^{\prime}=2$) as a transition from a nontrivial winding number of $C$ = +1 to a trivial winding number of $C$ = 0. This is due to the presence of a branch cut phase singularity where the transition point from nontrivial to trivial winding is where the gain enhancement is at a maximum (Fig. 4(b)). By contrast, changing only the real part $\varepsilon^{\prime}_{d}$ does not induce any phase transition as the complex reflection phase is completely continuous throughout the frequency range of interest (Fig. 4(c) and 4(d)). Thus, we see again that the loss-induced gain enhancement from the line defect is due to the topological protection of the gain resonance as inferred from the integer winding number as the topological invariant.
Figure 5: (a) Q factor of the quasi-BIC in the 1D system. At the defect loss value of 0.14 the resonance attains a maximum Q. (b) Q factor of the quasi-BIC in the 2D system.It should be noted that the y-axis of the Q factor vs defect loss plot is on a logarithmic scale.
To provide a clearer understanding of the nature of these loss-enhanced resonances in the 1D and 2D systems, we turn to examining the location of the resonance frequency in relation to the photonic bandstructures of the baseline periodic goemetries in 1D and 2D respectively. Focusing just on the real part of the bandstructure as the system is inherently non-Hermitian, the real part of the resonance frequency lies within the continuum in the periodic baseline. Doing the same calculation of the bandstructure in the reciprocal space of the 2D system, one can again see that the defect state lies within the continuum of states in the baseline bandstructure. Consequently, this shows that the loss-enhanced resonances observed in both systems have characteristics we associated with quasi-BIC’s due to their finite linewidth [34].
Examination of the field profiles and quality factors as measured by $\omega_{res}/FWHM$ where the denominator is the full width half max of the resonance peak further characterizes the behavior of the quasi-BIC’s. In the 1D case at the resonance peak in Figure 1(b), the field profiles of the TE field at a defect loss value of 0.14 is maximally localized at the center of the point defect. This creates a spatially symmetric mode profile that is protected by the underlying $C_{2}$ symmetry of the system. The quality factor $Q$ is also maximized at this value of loss as well at a very high value of about $4.5\times 10^{4}$ (Fig. 5(a)). This behavior can also be seen in the 2D system (Fig. 5(b)) as well the mode profile at the optimal loss value of 1.5 leads to a symmetry protected quasi-BIC about the center of the line defect. At this defect loss value, the $Q$ factor is also maximized as well reaching to a value $\sim 10^{5}$ (Fig. 5(b)).
@@
In summary, we have demonstrated that material loss, in contrast to the material’s real part of its permittivity can play the unexpected role of enhancing gain through a topological phase transition in 1D and 2D non-Hermitian photonic crystals. These resonances are additionally quasi-BIC’s whose $Q$ factors reach maximal values at the point of the phase transition. Our work highlights different the role that material loss and defects can play in non-Hermitian photonic structures. More broadly, it highlights how phase singularities can be used to engineer nonlinear photonic systems to have exceptional performance.
I Acknowledgments
@@
This material is based upon work supported by the National Science Foundation under grants No. ECCS-2146577 and CBET-2401080.

### v2

@@ II.2 Topological Phase Singularity Enhancement
Figure 5: (a) schematics of the baseline periodic and point-defected 2D non-Hermitian system. The baseline periodic system is shown on top with 12 unit cells of length a in the x-direction while infinitely periodic in the y-direction using Floquet periodic boundary conditions. Each unit cell is composed of air surrounding lossless rods of $\varepsilon_{a}=2$ and gain rods of $\varepsilon_{b}=2-0.1i$ with radii $0.17a$. The defected case is shown below which contains a lossy point defect composed of a rod with $\varepsilon_{d}$ and radius $r_{d}=0.3*a$. The defect is surrounded by 6 unit cells of the baseline periodic system on the left and right extents and made to be $C_{4}$ symmetric about the defect. (b) Gain vs frequency plots of the baseline periodic crystal compared against the lossy line defected case at loss values of 0.2, 0.32, and 0.35.
@@
The results observed thus far in 1D systems can also be extended to two-dimensional non-Hermitian system. We consider a baseline periodic 2D non-Hermitian photonic crystal with 12 unit cells that repeat in the $x$ direction of rods surrounded by air similar to the 1D baseline periodic structure (Fig. 5(a)). Within each unit cell, there is one lossless rod of $\varepsilon_{a}=2$ and a rod with gain of $\varepsilon=2-0.1i$. The rods have identical radii $0.17a$ where $a$ is the unit cell length. While the photonic crystal is finite in the x-direction, Floquet periodic boundary conditions are imposed to make the system infinitely periodic in the $y$-direction. In the defected case, a geometry similar to the 1D point-defected system is used where 6 unit cells surround either side of a lossy or lossless defect rod defined by a permittivity $\varepsilon_{d}=\varepsilon^{\prime}_{d}+\varepsilon^{\prime\prime}_{d}i$ and radius $r_{d}=0.3a$, resulting in $C_{4}$ rotational symmetry about the center of the point defect.
@@
Figure 6: (a), (b) Complex reflection phase as quantified by $Arg(r)$ vs $Im(\varepsilon_{d})$ and (c), (d) $Re(\varepsilon_{d})$ at 0 loss of the point defect respectively. Increasing $Im(\varepsilon_{d})$ leads to a transition from a winding number of 0 to +1, while the complex reflection phase remains continuous with changing $Re(\varepsilon_{d})$ in the absence of loss.
@@
As in the 1D system, increasing the loss in the point defect up to a critical value leads to multiple orders of magnitude gain enhancement (Fig. 5(b)) with extremely high quality factors due to the gain peaks corresponding to quasi-BICs (Figs. S3(b), S4(c), and S4(d)). This behavior can be understood by examining the integer winding of the complex reflection phase in the frequency vs $\varepsilon_{d}$ parameter space. Modulating the loss within the defect induces a topological phase transition around a critical loss value of $\varepsilon^{\prime\prime}_{d}=0.32$ (while maintaining $\varepsilon_{d}^{\prime}=2$) as a transition from a nontrivial winding number of $C$ = 0 to a trivial winding number of $C$ = +1 (Fig. 6(a)). This is due to the presence of a branch cut phase singularity where the transition point from nontrivial to trivial winding is where the gain enhancement is at a maximum (Fig. 6(b)). By contrast, changing only the real part $\varepsilon^{\prime}_{d}$ does not induce any phase transition as the complex reflection phase is completely continuous throughout the frequency range of interest (Fig. 6(c) and 4(d)). Thus, we observe again that the loss-induced gain enhancement from the line defect is due to the topological protection of the gain resonance as inferred from the integer winding number as the topological invariant.
@@
In summary, we have demonstrated that purely lossy defects can play the unexpected role of inducing both exceptional point and reflection pole singularities exhibiting nontrivial topology in non-Hermitian photonic crystals. In the case of reflection pole singularities, gain enhancement can be achieved and optimized through a topological phase transition in 1D and 2D non-Hermitian photonic crystals. Our work highlights the unique role that material loss and defects can play in non-Hermitian photonic structures. More broadly, it highlights further opportunities for engineering singularities to maximize performance in nonlinear photonic systems.
III Acknowledgments
@@
This material is based upon work supported by the National Science Foundation under grants No. ECCS-2146577. We thank Dr. Alexander Cerjan for helpful discussions related to this work.